OpenStudy (anonymous):

a 14-foot is learning aganist a building, with the base of the ladder 3 feet from the building. how high up on the building will the top of the ladder reach? ( round to the nearest tenth)

OpenStudy (anonymous):

\[14^2-3^2 = 196-9=187 \] now take the square root.